The base and altitude of a rectangle are in the ratio 3:2 and the area is 294 sq in. Find the dimensions of the rectangle.
Answer by lwsshak3(11628)

The base and altitude of a rectangle are in the ratio 3:2 and the area is 294 sq in. Find the dimensions of the rectangle.
***
b=base
a=altitude
b/a=3/2
b=3a/2
area=base*altitude
b*a=294
(3a/2)*a=294
3a^2=588
a^2=588/3
a^2=196
a=±√196=14
b=3a/2=42/2=21
dimensions of rectangle:
base=21
altitude=14